### v1

# Design the Quantum Instruction Set with the Cartan Coordinate Analysis Framework

Abstract
Most quantum compiling efforts rely on standard two-qubit basis gates, such as CX and iSWAP, to implement general quantum operations. However, with the advancement of quantum architecture design, more nonstandard two-qubit gates can now be implemented and calibrated on devices. Using these nonstandard gates may improve the performance of quantum computation. Despite this potential, it remains unclear how to efficiently incorporate these nonstandard gates into the quantum instruction set to enhance quantum advantage. To address this, we propose an analytical framework that facilitates the design of quantum instruction sets based on nonstandard gates. Our approach is grounded in the KAK decomposition and an analysis of the Cartan coordinate of two-qubit operations, enabling analytical conversion between any two two-qubit operations. This framework also demonstrates the lower and upper bounds of the conversion cost, revealing the relationship between the entangling power of two-qubit instructions and their Cartan coordinates. We further develop a compiler based on the analytical framework, which reduces the unitary decomposition cost with nonstandard two-qubit instructions. Using the proposed compiler, we evaluate various options for designing a quantum instruction set based on nonstandard gates. Our experiments demonstrate the efficiency of the proposed framework in the quantum instruction set design. Compared to the state-of-the-art method that is based on the numerical search, our framework reduces the time/resource overhead of exploring the instruction set design space by several orders of magnitude. Moreover, with the established framework, we propose feasible designs for the quantum instruction set by modeling real-world quantum processors, further promoting the quantum advantage.
I Introduction
With the development of quantum hardware, it becomes a feasible option to engineer quantum control pulses in a nonstandard way, rendering two-qubit gates not seen by the previous quantum instruction set architecture. Those gates, different from the well-known CX and iSWAP gates, are called ‘nonstandard’ gates. Existing works [1, 2, 3, 4] have demonstrated that using nonstandard two-qubit gates may improve the fidelity of near-term quantum circuits. Effective calibration schemes [1, 2, 4] for nonstandard two-qubit gates emerge at the same time, further supporting the usage of nonstandard gates in the quantum instruction set.
It is then natural to ask how to efficiently incorporate these nonstandard gates into the quantum instruction set to enhance quantum advantage. Unfortunately, existing tools cannot efficiently evaluate the options of the quantum instruction set design. Most existing works[1, 2] for the quantum instruction set design rely on numerical quantum compilers to evaluate the fidelity and latency of various benchmark programs decomposed with nonstandard gates. These numerical compilers will first construct a candidate circuit consisting of a series of nonstandard two-qubit gates with fixed gate parameters and single-qubit gates with tunable gate parameters, and then adjust the parameter of single-qubit gates such that the resulting circuit unitary is close to the unitary matrix of target quantum program. More fixed two-qubit gates and tunable single-qubit gates will be added to the candidate circuit if the approximation error is more than desired. Such a numerical search process will become slow if there are many tunable single-qubit gates in the candidate circuit. For example, with the state-of-the-art numerical compiler [1], decomposing the SWAP gate with the XX interaction gate $e^{i\frac{\pi}{28}X_{0}X_{1}}$ will take about two minutes, let alone to evaluate the performance of this gate in quantum applications whose circuit consists of several hundred of two-qubit subcircuits/operators. Peterson et al.[3] proposes to find the best candidate circuit closest to the target program in each search step analytically, largely reducing the compilation time. However, their work still does not overcome the path-like nature of the compilation process, and will inevitably incur higher compilation time if more search steps are needed for the target program.
To speed up the compilation process with nonstandard two-qubit gates thus assisting in the quantum instruction set design, we propose an analytical framework, named QCCA, based on the KAK decomposition, which enables an analytical conversion between any two two-qubit gates $g_{1}$ and $g_{2}$ if they are both not the SWAP gate. The conversion is based on the idea that a tunable single-qubit rotation gate can become a tunable two-qubit-level rotation gate when reflected by a fixed two-qubit gate, depending on the reflection used. Our proposed conversion can be finished in $10^{-3}$ seconds and in constant steps, which means the conversion time will not be affected by the pair of $g_{1}$ and $g_{2}$. With this tool, we can easily translate any quantum circuits written with the CX+U3 basis [5] or another commonly-used gate basis into the quantum gate basis consisting of the nonstandard two-qubit gate and single-qubit gates. The proposed framework also demonstrates the lower and upper bounds of the conversion cost, indicating the entangling power of nonstandard two-qubit gates.
We further propose an analytical compiler, which consists of a series of novel compiler optimizations for the conversion between nonstandard two-qubit gates. Our compiler greatly reduces the gate cost of using nonstandard gates in the instruction set to compile quantum programs, making its compilation results comparable to numerically optimal compilers. The proposed compiler is global and can finish the compilation within a fixed number of steps, having no local path-finding process existing in previous works, achieving thousands of times of speedup for the compilation. Based on the proposed compiler, we can compute and compare the effectiveness of nonstandard two-qubit gates of thousands of different gate configurations within seconds, significantly accelerating the quantum instruction set design.
Experimental results demonstrate that compared to the state-of-the-art method [1, 3, 2], our framework reduces the time/resource overhead of exploring the instruction set design space by several orders of magnitude. With the proposed framework, we evaluate various design options and propose feasible designs for the quantum instruction set by modeling real-world quantum processors, further promoting the quantum advantage.
II Background
In this section, we introduce the basic background knowledge of quantum computation and quantum compilation. For more details, we recommend [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19].
II-A Quantum Computation Basics
Qubit and quantum state: The basic information unit of quantum computation is called qubit. The state of a qubit is a unit vector $|\phi\rangle=\alpha|0\rangle+\beta|1\rangle,|\alpha|^{2}+|\beta|^{2}=1$ in the complex Hilbert space spanned by basis quantum states $\{|0\rangle,|1\rangle\}$.
Single-qubit gates: The state of a qubit can be manipulated by single-qubit quantum gates/operators, e.g., Pauli operators
$Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},Y=\begin{pmatrix}0&-i\\ i&0\end{pmatrix},I=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},$
and rotation gates $Z(\beta)=e^{i\beta Z}=\cos\beta I+i\sin\theta Z,X(\beta)=e^{i\beta X},Y(\beta)=e^{i\beta Y}$. Other commonly used gates include the phase gate $S=\begin{pmatrix}1&0\\ 0&i\end{pmatrix}$ and the Hadamard gate $H=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$.
Entangling quantum gates: Likewise, there are n-qubit gates that aim to manipulate the quantum state of n qubits simultaneously, such as the two-qubit gate $CX=\begin{pmatrix}I&O\\ O&X\end{pmatrix}$. N-qubit quantum gates are essentially $2^{n}\times 2^{n}$ unitary matrices. An n-qubit entangling gate can transform some product states of n qubits into entangled states that cannot be expressed as the tensor product of single-qubit quantum states. For more details, please refer to [6].
To differentiate gates applied to different qubits, this work uses subscripts, e.g., $Z_{0}$ means to apply the Z gate on qubit $q_{0}$, $H_{0,1}$ means to apply H gates on both $q_{0}$ and $q_{1}$.
II-B Cartan Decomposition and Local Equivalence
N-qubit quantum operators can be decomposed into the product of a series of single-qubit gates and two-qubit gates, i.e., a quantum circuit, with Cartan decomposition [7]. Moreover, any two-qubit unitary matrix $U\in SU(4)$ can be decomposed into the following canonical form (aka KAK decomposition):
$U=k_{g}(a\otimes b)e^{i(\eta_{x}X_{0}X_{1}+\eta_{y}Y_{0}Y_{1}+\eta_{z}Z_{0}Z_{1})}(c\otimes d),$ (II.1)
where the global phase $k_{g}\in\{1,i\}$, $a,b,c,d\in SU(2)$, $0\leq|\eta_{z}|\leq\eta_{y}\leq\eta_{x}\leq\frac{\pi}{4}$. If $\eta_{x}=\frac{\pi}{4}$, $\eta_{z}\geq 0$. $(\eta_{z},\eta_{y},\eta_{x})$ are called Cartan coordinate of $U$. Two-qubit unitary matrices $U_{1}$ and $U_{2}$ are called locally equivalent if they have the same Cartan coordinate. Based on the notion of local equivalence, any $SU(4)$ unitary can be categorized into the following three templates:
$D_{a}(\theta_{x})$ $=e^{i\theta_{x}X_{0}X_{1}},$ (II.2)
$D_{b}(\theta_{x},\theta_{y})$ $=e^{i(\theta_{x}X_{0}X_{1}+\theta_{y}Y_{0}Y_{1})}$ (II.3)
$D_{c}(\theta_{x},\theta_{y},\theta_{z})$ $=e^{i(\theta_{x}X_{0}X_{1}+\theta_{y}Y_{0}Y_{1}+\theta_{z}Z_{0}Z_{1})}$ (II.4)
With Cartan decomposition, it is possible to decompose any n-qubit quantum operator with a fixed gate group, such as the group consisting of the CX gate and all single-qubit gates. Such a gate group are also called a quantum instruction set. Gates in the instruction set are called basis gates. For a two-qubit basis gate $g$, and a quantum program $P$, if we need $n$ invocations of $g$ to implement $P$, we define the decomposition cost of $P$ with $g$ is $n$.
Without loss of generality, this work only considers the three templates above for the quantum instruction set design. The template $D_{a}$ is widely available in various quantum computing platforms, including IBM and Rigetti superconducting devices [5, 20]. Chen et al. [4] recently proposes a pulse scheme to implement the template $D_{b}$ on superconducting devices. Google’s Sycamore processor [21] supports the fsim gate, which is locally equivalent to $D_{b}$.
III Theory of Cartan Coordinate Transformation
Figuring out how to analytically convert between two-qubit gates is important for the quantum instruction set design since now only a few analytical computation is sufficient to reason about the overhead of a specific two-qubit gate in implementing quantum applications. Such an efficient reasoning is not possible if we use the computationally expensive numerical search to find the decomposition of quantum applications.
The proposed theory of Cartan coordinate transformation studies how to analytically implement other two-qubit gates with an arbitrary two-qubit (entangling) gate. The proposed theory is motivated by the Solovay-Kitaev theorem [10] where various rotation gates can be obtained by perform T gate on the rotated Bloch sphere. Likewise, for the two-qubit case, it may be possible to implement an arbitrary two-qubit quantum operation by repeating the process: first rotate the 2q Hilbert space with the two-qubit basis gate, then perform some single-qubit rotation gates, and then rotate the Hilbert space back. Indeed, the circuit $e^{i\beta Z_{1}Z_{2}}=CX\cdot Z(\beta)\cdot CX$ exactly follows this idea. Now we generalize this idea to an arbitrary two-qubit (entangling) gate:
Proposition III.1.
For a fixed $\theta_{x}$, $\frac{\pi}{4}\geq\theta_{x}>0$,
$e^{i\eta X_{0}X_{1}}=Z_{1}(\tau)D_{a}Z_{1}(\beta)D_{a}^{\dagger}Z_{1}(\tau)(-iZ_{1}),$ (III.1)
where $\eta,\tau$ satisfy
[…]
Without ambiguity, we call $Z_{1}(\tau)$, $Z_{1}(\tau)(-iZ_{1})$ calibration operators of $D_{a}Z_{1}(\beta)D_{a}^{\dagger}$, making the latter becoming the XX-rotation operator.
Proposition III.2.
For a fixed $\theta_{x},\theta_{y}$, $\frac{\pi}{4}\geq\theta_{x}\geq\theta_{y}>0$,
[…]
where $\eta_{x},\eta_{y},\tau_{0},\tau_{1}$ satisfy
[…]
Remark.
Proposition III.1 and III.2 can be proved with direct computation. From the perspective of Lie algebra theory, the right hand side (RHS) term of Equation III.1 is in the space $exp(V)$, where $V$ is the Cartan algebra $i\{X_{0}X_{1},X_{0}Y_{1},I_{0}Z_{1}\}$. Then according to the Cartan decomposition theorem [7], $exp(V)=exp(i\{I_{0}Z_{1}\})exp(i\{X_{0}X_{1}\})exp(i\{I_{0}Z_{1}\})$. To some extent, this formula is just the restatement of Equation III.1 since $exp(i\{IZ\})$ is exactly the Z-rotation gate on $q_{1}$. We can explain Equation III.3 from the perspective of Lie algebra theory in the same manner.
Remark.
From the above angle relations, it is easy to see that $\frac{\pi}{2}-\theta_{x}$ has a similar effect as $\theta_{x}$, making it okay to assume $\frac{\pi}{4}\geq\theta_{x}>0$. Moreover, for $D_{a},D_{b}$, $D_{a}^{\dagger}$ and $D_{b}^{\dagger}$ can always be implemented with $ZD_{a}Z(YD_{a}Y),ZD_{b}Z$.
Proposition III.1 and III.2 reveal how to transform a 2q gate with a fixed Cartan coordinate into a 2q gate with parameterized Cartan coordinate. Those two propositions can be used to generate formulas for other transformations based on $D_{a},D_{b},D_{c}$, with the following corollary:
Corollary III.2.1.
We observe the following rules for parameterized Cartan coordinate transformation:
  1. 1.
    (Pauli conversion) By multiplying $H_{1}S_{1}H_{1}H_{0}S_{0}H_{0}$ and $H_{1}S^{\dagger}_{1}H_{1}H_{0}S^{\dagger}_{0}H_{0}$ to the left and right side of the LHS and RHS term respectively in Equation III.1, we see $D_{a}Y(\beta)D_{a}^{\dagger}$ is locally equivalent to $e^{i\eta X_{0}X_{1}}$ up to some local Y-rotation gates.
  2. 2.
    (Reduce) $D_{a}Z(\beta)D_{a}=D_{a}Z(\beta-\frac{\pi}{2})D_{a}^{\dagger}Z(\frac{\pi}{2})$, $D_{a}Y(\beta)\\ D_{a}=D_{a}Y(\beta-\frac{\pi}{2})D_{a}^{\dagger}Y(\frac{\pi}{2})$, $D_{b}Z(\beta)D_{b}=D_{b}Z(\beta-\frac{\pi}{2})D_{b}^{\dagger}Z(\frac{\pi}{2})$;
  3. 3.
    (Downgrade I) $D_{b}Y(\theta)D_{b}^{\dagger}=D_{a}Y(\theta)D_{a}^{\dagger}$ (assuming the same $\theta_{x}$), $D_{c}Z(\theta)D_{c}^{\dagger}=D_{b}Z(\theta)D_{b}^{\dagger}$;
  4. 4.
    (Downgrade II) $D_{b}Y(\theta)D_{b}=e^{i2\theta_{y}Y_{0}Y_{1}}D_{a}Y(\theta)D_{a}$, and then Rule 2; If $D_{c}^{\dagger}$ is not available, $D_{c}Z(\beta)D_{c}=e^{i2\theta_{z}Z_{0}Z_{1}}D_{b}Z(\theta)D_{b}$, and then reuse Rule 2.
Note that, since $e^{i2\theta_{y}Y_{0}Y_{1}}$ is commutable with Y-rotation gates, $e^{i2\theta_{y}Y_{0}Y_{1}}D_{a}Y(\theta)D_{a}$ is locally equivalent to $e^{i2\theta_{y}Y_{0}Y_{1}}e^{i\eta X_{0}X_{1}}$ by Proposition III.1. The case for $e^{i2\theta_{z}Z_{0}Z_{1}}D_{b}Z(\theta)D_{b}$ is similar.
Takeaway I We can use any two-qubit entangling gate to analytically achieve any desired Cartan coordinate.
Roughly speaking, with Proposition III.1, Proposition III.2 and Corollary III.2.1, we may think six invocations of $D_{a}$ or $D_{b}$ or $D_{c}$ are enough to achieve any two-qubit quantum gate. Unfortunately, this is theoretically not possible.
Firstly, it is obvious that in Proposition III.1, $\eta\in[-2\theta_{x},2\theta_{x}]$; in Proposition III.2, $|\eta_{x}\pm\eta_{y}|\leq 2(\theta_{x}\pm\theta_{y}$). It means, when $\theta_{x}$ is small, we cannot guarantee $\eta\in[-\frac{\pi}{4},\frac{\pi}{4}]$, which is a basic requirement by the universal KAK decomposition.
Secondly, while the construction may affect the required number of invocations of the specific two-qubit gate, the following theorem further tells us that, regardless of the construction used, it is not possible to use only six invocations (of any two-qubit entangling gate) to achieve any 2q unitary matrix:
Theorem III.3 (Lower Bound of the Decomposition Cost).
Assume the count of $D_{a}$ in the decomposed circuit of $U\in SU(4)$ is $N$. If $N\theta_{x}<\frac{\pi}{4}$, then $sin(k_{t}(U))\leq\sin N\theta_{x}$, where $k_{t}(U)$ is the L1 norm of the Cartan coordinate of $U$.
Proof.
Please see the appendix. ∎
Corollary III.3.1.
Similarly, the following conclusions can be obtained by repeatedly applying Theorem III.3:
  1. 1.
    For $D_{b}$, if $N(\theta_{x}+\theta_{y})<\frac{\pi}{4}$ ($N$ is the decomposition cost of $U$ with $D_{b}$), then $\sin\,k_{t}(U)\leq\sin N(\theta_{x}+\theta_{y})$;
  2. 2.
    For $D_{c}$, if $N(\theta_{x}+\theta_{y}+|\theta_{z}|)<\frac{\pi}{4}$ ($N$ is the decomposition cost of $U$ with $D_{c}$), then $\sin\,k_{t}(U)\leq\sin N(\theta_{x}+\theta_{y}+|\theta_{z}|)$.
Takeaway II Using small rotation angles for $D_{a},D_{b}$ or $D_{c}$ will lead to larger worst-case decomposition cost for general quantum operations, thus $D_{a},D_{b}$ or $D_{c}$ being of lower entangling power.
With Theorem III.3, it is easy to see that exactly $k$ invocations of $CX^{1/k}$ is needed to achieve the CX gate. If the gate duration or error rate of $CX^{1/k}$ is no more than $1/k$ of those of the CX, $CX^{1/k}$ will be a perfect drop-in replacement for the CX gate in quantum applications. However, due to the physical limitation, this assumption will be invalidated as $k$ becomes larger, requiring a careful design for the value $k$. On the other hand, considering $CX^{\alpha},(\alpha>1)$ may not be a good option since we have shown that $CX^{\alpha}$ is no difference from $CX^{\alpha\%1}$ regarding the KAK decomposition, let alone a large $\alpha$ often leads to a higher gate latency and error rate.
Overall, to design an efficient quantum instruction set, we need to balance various factors, such as the decomposition cost, the gate latency, and the gate fidelity.
IV Compilation with Arbitrary 2Q Gate
Section III mostly discusses the cost of turning a fixed Cartan coordinate into a parameterized one. To further evaluate the efficiency of a specific 2q gate, we need to examine its functioning in general quantum circuits. In this section, we propose a compiler which analytically compiles general quantum programs with an arbitrary 2q entangling gate. Theoretically, there is unlimited ways to apply Proposition III.1, III.2 and Corollary III.2.1 during the compilation process. In our compiler design, we choose to utilize those constructions globally, that is, once the target unitary is given, our compiler immediately determines the required number of basis 2q gates, and directly produces the compiled circuit without using the path-like searching in state-of-the-art works [1, 3, 2].
IV-A Case I: with $D_{a}$
We consider the compilation with $D_{a}$ as the only two-qubit gate in the instruction set. To begin with, we consider the case where $\frac{\pi}{4}\geq\theta_{x}\geq\frac{\pi}{8}$. Then with the construction in Proposition III.1, any Cartan coordinate can be achieved by:
$e^{i(\sum_{\sigma\in[X,Y,Z]}\eta_{\sigma}\sigma\otimes\sigma)}=e^{i\eta_{y}Y_{0}Y_{1}}e^{i\eta_{x}X_{0}X_{1}}e^{i\eta_{z}Z_{0}Z_{1}}=$ (IV.1)
$\quad S_{0}S_{1}(Z_{0}(\tau_{y})D_{a}Z_{0}(\beta_{y})D_{a}^{\dagger}Z_{0}(\tau_{y})(-iZ_{0}))S_{0}^{\dagger}S_{1}^{\dagger}\cdot$
$\quad(Z_{1}(\tau_{x})D_{a}Z_{1}(\beta_{x})D_{a}^{\dagger}Z_{1}(\tau_{x})(-iZ_{1}))\cdot$
$\quad H_{0}H_{1}(Z_{0}(\tau_{z})D_{a}Z_{0}(\beta_{z})D_{a}^{\dagger}Z_{0}(\tau_{z})(-iZ_{0}))H_{0}H_{1}$
(Optimization I) According to Equation IV.1, if $\tau_{y}=-\tau_{x}=\pm\frac{\pi}{4}$ ($\Longleftrightarrow\theta_{x}=\frac{\pi}{4}$), $Z_{0}(\tau_{y})(-iZ_{0}))S_{0}^{\dagger}S_{1}^{\dagger}Z_{1}(\tau_{x})=-iZ_{0}Z_{1}$ or $-iI$, thus commuting with $D_{a}^{\dagger}$, indicating that the $D_{a}^{\dagger}$ in $e^{i\eta_{y}Y_{0}Y_{1}}$ can cancel with the $D_{a}$ in $e^{i\eta_{x}X_{0}X_{1}}$. Similarly, with gate exchanging between $D_{a}$, $D_{a}^{\dagger}$ and X-rotation gates, we can extract $u=D_{a}^{\dagger}(-iZ_{1})H_{0}H_{1}D_{a}$ from the subcircuit for $e^{i\eta_{x}X_{0}X_{1}}e^{i\eta_{z}Z_{0}Z_{1}}$. If $\theta_{x}=\frac{\pi}{4}$, then $u=(-iZ_{1})H_{0}H_{1}\cdot SWAP\cdot S_{0}S_{1}D_{a}^{\dagger}S_{0}^{\dagger}S_{1}^{\dagger}$ only requires one invocations of $D_{a}$. Therefore, if $\theta_{x}=\frac{\pi}{4}$, only 3 invocations of $D_{a}$ is needed for the general KAK decomposition. Indeed, if $\theta_{x}=\frac{\pi}{4}$, $D_{a}$ is locally equivalent to the CX gate.
(Optimization II) For $u=e^{i\eta X_{0}X_{1}}$, let $m=\lceil\frac{\eta}{\theta_{x}}\rceil$, then we can implement $u$ with $m$ invocations of $D_{a}$. We can construct the decomposition circuit as follows: since $m\theta_{x}\geq\eta$, then $\eta-(m-2)\theta_{x}\leq 2\theta_{x}$, i.e., $u(D_{a}^{\dagger})^{m-2}$ can be implemented with only 2 invocations of $D_{a}$, according to Proposition III.1. If $m=1$, then 2 invocations of $D_{a}$ is sufficient for $u$ (if further $\eta=\theta_{x}$, then only one invocation is needed). According to Theorem III.3, this optimization leads to the optimal compilation of $e^{i\eta X_{0}X_{1}}$ with $D_{a}$.
Combining with the above optimizations, there are two ways for decomposing 2q unitary operators with $D_{a}$: 1) indirect synthesis: using $D_{a}$ to implement CX, then using CX for achieving general 2q gates; 2) direct synthesis: directly using $D_{a}$ to achieve any Cartan coordinate. The cost of CX-based synthesis for achieving the Cartan coordinate $(\eta_{x},\eta_{y},\eta_{z})$ is $3k$, where $k$ is the number of invocations of $D_{a}$ required to implement CX. The cost of direct synthesis for the same Cartan coordinate is $M=\lceil\frac{\eta_{x}}{\theta_{x}}\rceil+\lceil\frac{\eta_{y}}{\theta_{x}}\rceil+\lceil\frac{|\eta_{z}|}{\theta_{x}}\rceil+C$ ($C$ is the compensating term determined by corner cases of Optimization II). Combing these two synthesis methods, to implement any 2q unitary gate $U$, at most $N_{U}$ invocations of $D_{a}$ is needed:
$N_{U}=\begin{cases}\min(3\frac{\pi}{4\theta_{x}},M),\ if\ \frac{\pi}{4\theta_{x}}\in\mathbb{Z}\\ M.\end{cases}$ (IV.2)
When $D_{a}$ is more efficient than CX? We first consider the case when $\frac{\pi}{4\theta_{x}}\notin\mathbb{Z}$. For the implementation of the 2q operation $u=e^{i\eta_{x}X_{0}X_{1}}$ ($\eta_{x}\leq 2\theta_{x}\leq\frac{\pi}{4}$), it requires two invocations of $D_{a}$ or two invocations of CX. Taking the execution time as the metric, then for this $u$, $D_{a}$ is more advantageous than CX since $D_{a}$ is faster to prepare than the CX gate. However, for 2q operations $e^{i\eta_{x}X_{0}X_{1}}e^{i\eta_{y}Y_{0}Y_{1}}$ or $e^{i\eta_{x}X_{0}X_{1}}e^{i\eta_{y}Y_{0}Y_{1}}e^{i\eta_{z}Z_{0}Z_{1}}$ ($\eta_{x}+\eta_{y}+|\eta_{z}|<k\theta_{x}$), $D_{a}$ is more efficient than CX unless the time of $D_{a}$ is no more than $3/k$ of the time of CX since for those 2q operations the decomposition cost with $D_{a}$ is about $\frac{k}{3}\times$ that with CX. The discussion over the case where $\frac{\pi}{4\theta_{x}}\in\mathbb{Z}$ is similar.
Takeaway III Compared to the CX gate, using the basis 2q gate with a small rotation angle may reduce the latency and improve the fidelity of executing the unitary operator with a small Cartan coordinate.
IV-B Case II: with $D_{b}$
We then consider the compilation with $D_{a}$ as the only two-qubit gate in the instruction set. According to the Downgrade Rule I in Corollary III.2.1, the basic compiling flow with $D_{b}$ follows the one with $D_{a}$, except that $D_{b}$ allows more optimization opportunities.
(Optimization I) For the operation $u=e^{i(\eta_{x}X_{0}X_{1}+\eta_{y}Y_{0}Y_{1})}$, with Proposition III.1 and Corollary III.2.1,
$u$ $=a_{1}^{Y}D_{b}Y_{1}(\beta_{x})D_{b}^{\dagger}b_{1}^{Y}a_{0}^{X}D_{b}X_{0}(\beta_{y})D_{b}^{\dagger}b_{0}^{X}$ (IV.3)
$=a_{1}^{Y}a_{0}^{X}D_{b}Y_{1}(\beta_{x})X_{0}(\beta_{y})D_{b}^{\dagger}b_{1}^{Y}b_{0}^{X},$ (IV.4)
where $a_{1}^{Y},b_{1}^{Y}$ are Y-rotation-based calibration operators on $q_{1}$, $a_{0}^{X},b_{0}^{X}$ are X-rotation-based calibration operators on $q_{0}$. Equation IV.6 is due to the fact that $b_{1}^{Y}$ commutes with $D_{b}X_{0}(\beta_{y})D_{b}^{\dagger}=e^{i\theta_{y}Y_{0}Y_{1}}X_{0}(\beta_{y})e^{-i\theta_{y}Y_{0}Y_{1}}$, and likewise $a_{0}^{X}$ commutes with $D_{b}Y_{1}(\beta_{x})D_{b}^{\dagger}$.
(Optimization II) If $\theta_{x}=\frac{\pi}{4}$, then for $u=e^{i(\eta_{x}X_{0}X_{1}+\eta_{z}Z_{0}Z_{1})}$, with Proposition III.1 and Corollary III.2.1, we have
$u$ $=a_{1}^{Y}D_{b}Y_{1}(\beta_{x})D_{b}^{\dagger}b_{1}^{Y}H_{0,1}a_{0}^{Y}D_{b}Y_{0}(\beta_{z})D_{b}^{\dagger}b_{0}^{Y}H_{0,1}$ (IV.5)
$=a_{1}^{Y}a_{0}^{X}D_{b}Y_{1}(\beta_{x})Y_{0}(\beta_{z})D_{b}^{\dagger}b_{1}^{Y}H_{0}H_{1},$ (IV.6)
This is because $b_{1}^{Y}=H_{1},a_{0}^{Y}=H_{0}$, if $\theta_{x}=\frac{\pi}{4}$.
(Optimization III) If $\theta_{x}=\frac{\pi}{4},\theta_{y}=\frac{\pi}{8}$, $u=D_{b}Y_{0}(\beta_{0})Z_{1}(\beta_{1})\\ Y_{1}(\beta_{2})Z_{1}(\beta_{1})D_{b}$ is locally equivalent to any 2q unitary by adjusting $\beta_{0},\beta_{1},\beta_{2}$.
(Optimization IV) For $u=e^{i\eta_{x}X_{0}X_{1}}$, $m=\lceil\frac{\eta_{x}}{\theta_{x}}\rceil$, then $m^{\prime}=2\lceil\frac{m}{2}\rceil$ invocations of $D_{b}$ is sufficient for implementing $u$. Indeed, since $\eta_{x}\leq m^{\prime}\theta_{x}$, then $\frac{\eta_{x}}{m^{\prime}/2}\leq 2\theta_{x}$, thus with Proposition III.1, $2*(m^{\prime}/2)=m^{\prime}$ invocations of $D_{b}$ is sufficient for $u$, based on the construction $D_{b}Y_{1}(\beta)D_{b}^{\dagger}$.
(Optimization V) For $u=e^{i(\eta_{x}X_{0}X_{1}+\eta_{y}Y_{0}Y_{1})}$ ($\eta_{x}\geq\eta_{y}$), let $m=\lceil\frac{\eta_{x}}{\theta_{x}}\rceil$, $n=\lceil\frac{\eta_{y}}{\theta_{y}}\rceil$. If $m\geq n$, then $m$ invocations of $D_{b}$ is sufficient for implementing $u$. Otherwise, let $n^{\prime}=\lceil\frac{\eta_{y}-m\theta_{y}}{\theta_{x}}\rceil$, then $m+2\lceil\frac{n^{\prime}}{2}\rceil$ $D_{b}$ is sufficient. Here the corner case $m=1$ is ignored for simplicity.
For the former case, the decomposition circuit can be constructed as follows: let $D_{b^{\prime}}=XD_{b}X$, then $u^{\prime}=u(D_{b^{\prime}}^{\lceil\frac{m-n-1}{2}\rceil})^{\dagger}(D_{b}^{n-1+\lfloor\frac{m-n-1}{2}\rfloor})^{\dagger}$, its Cartan coordinate satisfies $\eta_{x}^{\prime}\leq 2\theta_{x},\eta_{y}^{\prime}\leq 2\theta_{y}$, thus $u^{\prime}$ (or $u^{\prime}$) can be implemented with only 2 (resp. m) $D_{b}$ with Optimization I. For the latter case, for $u^{\prime}=(D_{b}^{\dagger})^{m-2}u$, one usage of $D_{b}Y_{1}(\beta)D_{b}^{\dagger}$ is sufficient to cancel out the $\eta_{x}^{\prime}$ of $u^{\prime}$. Then, the remaining $\eta_{y}^{\prime\prime}=\eta_{x}-m\theta_{y}$, and we can reuse Optimization IV to construct the circuit that cancels out the remaining $e^{i\eta_{y}^{\prime\prime}Y_{0}Y_{1}}$.
(Optimization VI) For $u=e^{i(\eta_{x}X_{0}X_{1}+\eta_{y}Y_{0}Y_{1}+\eta_{z}Z_{0}Z_{1})}$ ($\eta_{x}\geq\eta_{y}\geq\eta_{z}$), let $m=\lceil\frac{\eta_{x}}{\theta_{x}}\rceil$, $n=\lceil\frac{\eta_{y}}{\theta_{y}}\rceil$. If $m\geq n$, let $l=\lceil\frac{|\eta_{z}|}{\theta_{x}}\rceil$, then $m+2\lceil\frac{l}{2}\rceil$ invocations of $D_{b}$ is sufficient for implementing $u$; otherwise, let $\eta_{x}^{\prime}=\max(\eta_{y}-m*\theta_{y},|\eta_{z}|),\eta_{y}^{\prime}=\min(\eta_{y}-m*\theta_{y},|\eta_{z}|)$, $m^{\prime}=\lceil\frac{\eta^{\prime}_{x}}{\theta_{x}}\rceil$, $n^{\prime}=\lceil\frac{\eta^{\prime}_{y}}{\theta_{y}}\rceil$, if $m^{\prime}\geq n^{\prime}$, then overall $m+m^{\prime}$ invocations of $D_{b}$ is sufficient; otherwise, let $l=\lceil\frac{\eta_{y}^{\prime}-m^{\prime}\theta_{y}}{\theta_{x}}\rceil$, then $m+m^{\prime}+2\lceil\frac{l}{2}\rceil$ invocations of $D_{b}$ should be enough. Here corner cases like $m=1$ or $m^{\prime}=1$ are ignored for simplicity.
The construction of the decomposed circuit for Optimization VI is similar to that for Optimization V, and we omit it here for simplicity.
Overall, like the compiling flow with $D_{a}$, there are two ways to compile unitary operators with $D_{b}$: direct synthesis and indirect synthesis. The indirect synthesis will first implement special gates (e.g., CX with Optimization II, B gate with Optimization III) and then use those special gates to do the compilation. The direct synthesis follows the idea of direct synthesis with $D_{a}$. The compilation with $D_{b}$ will first compute the decomposition cost of each synthesis scheme and pick the one producing the lowest decomposition cost to do the actual compilation. Comparing to $D_{a}$, $D_{b}$ tends to have lower decomposition cost (if $\theta_{x}$ is the same), since we have much more optimization opportunities with $D_{b}$.
IV-C Case III: with $D_{c}$
Finally, we consider the compilation with $D_{c}$ as the only two-qubit gate in the instruction set. There are generally two ways to compile 2q quantum unitary operators with $D_{c}$:
The first way is indirect synthesis, which considers the identities
$D_{c}X_{0}D_{c}X_{0}=e^{i2\theta_{x}X_{0}X_{1}},\quad$ $D_{c}Y_{0}D_{c}Y_{0}=e^{i2\theta_{y}Y_{0}Y_{1}},$ (IV.7)
$D_{c}Z_{0}D_{c}Z_{0}$ $=e^{i2\theta_{z}Z_{0}Z_{1}}.$
Then we can follow the flow of compiling with $D_{a}$. As for the optimal rotation angle of $D_{a}$, considering the periodic property of n-qubit Pauli rotations, we can set the optimal rotation angle of $D_{a}$ as
[…]
Comparing to $D_{a}$ and $D_{b}$, this way of using $D_{c}$ will lead to about 2x decomposition cost for $e^{i\eta X_{0}X_{1}}$, $\eta\in[-\theta_{x}^{\prime},\theta_{x}^{\prime}]$.
The second way is to solve the following Equation that is based on Proposition III.2,
$(\eta_{x},\eta_{y},\eta_{z})=\sum_{i}^{n}(k^{x}_{i},k^{y}_{i},k^{z}_{i})P_{i}$ (IV.9)
Where $P_{i}$ is the permutation matrix of three dimensional tuples, the Cartan coordinate $(k^{x}_{i},k^{y}_{i},k^{z}_{i})$ is either from $D_{c}Z_{1}(\beta)D_{c}$, $D_{c}X_{1}(\beta)D_{c}$ or $D_{c}Y_{1}(\beta)D_{c}$, the decomposition cost $n$ is bounded by $[(\eta_{x}+\eta_{y}+|\eta_{z}|+m\frac{\pi}{2})/\theta_{x},(\eta_{x}+\eta_{y}+|\eta_{z}|+m\frac{\pi}{2})/|\theta_{z}|]$, $m\in\mathbb{Z}$. Unfortunately, it is not a trivial task to solve the Equation IV.9 analytically. A feasible way for solving Equation IV.9 is to use the numerical search-based method as in [1]. Since this work focuses on the analytical compilation, we leave the numerical solver design as a future work.
On the other hand, if we design the rotation angles in $D_{c}$ carefully, we may solve Equation IV.9 analytically in certain cases. For example, for the SWAP gate, if $\frac{\pi}{4\theta_{y}}\in 4\mathbb{Z},\theta_{x}+|\theta_{z}|\geq 2\theta_{y}$, we can simply set $P_{2k+1}=I$, $P_{2k+2}=\sigma(1,3)$, $(k^{x}_{i},k^{y}_{i},k^{z}_{i})$ to be by $D_{c}Y_{1}(\beta)D_{c}$ and $k^{x}_{i}+k^{z}_{i}=2k^{y}_{i}$. In this case, the decomposition cost for the SWAP gate with $D_{c}$ is just $\frac{\pi}{4\theta_{y}}$.
IV-D Compiling General Quantum Circuits
Now, we can compile any n-qubit quantum unitary gate with $D_{a},D_{b}$ or $D_{c}$. We can first decompose the n-qubit quantum unitary gate into a group of 2q and 1q gates and then use Algorithm 1 to compile any 2q unitary gates.
Input: operator to be decomposed $U_{y}$, basis gate $U_{x}$;
Do KAK decomposition on $U_{x}$, and get $U_{x}=(a_{x}\otimes b_{x})D_{x}(c_{x}\otimes d_{x})$;
Do KAK decomposition on $U_{y}$, and get $U_{y}=(a_{y}\otimes b_{y})D_{y}(c_{y}\otimes d_{y})$;
if $D_{x}$ is of the $D_{a}$ type then
       $C$ = Compiling $D_{y}$ as in Section IV-A;
else if $D_{x}$ is of the $D_{b}$ type then
       $C$ = Compiling $D_{y}$ as in Section IV-B;
else if $D_{x}$ is of the $D_{c}$ type then
       $C$ = Compiling $D_{y}$ as in Section IV-C;
$C^{\prime}$ = Replacing $D_{x}$ in $C$ with $(a^{\dagger}_{x}\otimes b^{\dagger}_{x})U_{x}(c^{\dagger}_{x}\otimes d^{\dagger}_{x})$;
Output: $U_{y}=(a_{y}\otimes b_{y})C^{\prime}(c_{y}\otimes d_{y})$;
Algorithm 1 Decomposition with any 2Q Gate
V Evaluation
In this section, we first compare the proposed analytical compilation framework with the state-of-the-art compiler that can also work with nonstandard two-qubit basis gates. We then evaluate various options of the quantum instruction set design with the proposed analytical compiler. For the calibration of nonstandard gates, we follow existing schemes [2, 4].
V-A Experiment Setting
Platform: All experiments are performed with Python 3.8.10, on a Ubuntu 20.04 server with 20 CPU cores and 32GB RAM.
Baseline: We first compare the proposed analytical compiler QCCA to the numerical search-based compiler NuOp [1, 2], which also supports compilation with nonstandard two-qubit gates. Since NuOp is designed to compile two-qubit unitary blocks in the general quantum circuits, we focus on compiling two-qubit quantum operations when comparing to NuOp.
We then compare QCCA to the XXDecomposer [3] in Qiskit, which can also works with nonstandard two-qubit basis gates, but only for those of the $D_{a}$ type. The basic idea of XXDecomposer is to analytically compute the circuit which has the nearest Cartan coordinate to the target Cartan coordinate, within the circuit subspace $P_{n}$ which consists of $n$ $D_{a}$ gates. If the target Cartan coordinate is not in $P_{n}$, then XXDecomposer will consider $P_{n+1}$ until the compiled circuit of the target Cartan coordinate is found.
Metrics: The metrics considered is the compilation time for quantum operations, and the decomposition cost, which means the required number of the basis two-qubit gate to implement a specific quantum unitary. We also consider the fidelity and the latency of the decomposed circuit. The circuit fidelity is computed with the formula $\prod_{g\in circ}(1-e_{g})$, where $e_{g}$ is the error rate of the gate $g$. The circuit latency is normalized to the count of CX gates.
(a) For QCCA with $D_{a}$
(b) For QCCA with $D_{b}$
(c) For QCCA with $D_{c}$
Figure 1: Comparing the compilation time by QCCA to that by NuOp. Both NuOp and QCCA adopt the same 2q basis gate in compilation. The y axis is $\log_{10}\frac{t_{NuOp}}{t_{QCCA}}$, where $t_{NuOp}$ and $t_{QCCA}$ are the compilation time by NuOp and QCCA, respectively. As for the X axis, for (a), x axis value $k$ means to use $D_{a}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k}$; for (b), x axis value $k$ means to use $D_{b}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k},\theta_{y}=\frac{\pi}{8k}$; for (c), x axis value $k$ means to use $D_{c}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k},\theta_{y}=\frac{\pi}{8k},\theta_{z}=\frac{\pi}{16k}$.
Benchmark: For the comparison to the baseline, we use the benchmark consisting of the CX gate, the SWAP gate, and 100 random unitary operators. The CX gate is the mostly widely used two-qubit gates in various existing quantum circuits, while the SWAP gate is essential for executing quantum circuits with limited qubit connectivity. The random unitary can be used to demonstrate the average decomposition cost of compilers. Due to the randomness of the numerical compilation in the baseline, we will repeat the compilation of each benchmarking 2q unitary by 100 times and compute the averaged metrics.
We design the quantum instruction set in an application-aware manner. Thus, in this part of evaluation, we consider four benchmark programs: Quantum Approximate Optimization Algorithm (QAOA), Quantum Fourier transform (QFT), Bernstein-Vazirani algorithm (BV) and Pauli Evolution programs (PauliEvo).
QAOA is a promising quantum application in the NISQ (Noisy Intermediate-Scale Quantum) era, designed to solve combinatorial optimization problems. QAOA programs have also been widely used to benchmarking the performance of quantum processors [22]. For QAOA, Our experiments use the MAXCUT ansatz over 100 random regular graphs where we will have a ZZ interaction between two nodes with a probability 0.3.
QFT is a key subroutine in many important quantum applications, such as Shor’s algorithm [23] and the quantum adder [24]. The QFT program with $n$ qubits contains $n(n-1)/2$ controlled-rotation gates, with rotation angles as small as $2\pi/2^{n}$.
BV algorithm is used to learn a string encoded in a oracle function and is important for quantum cryptography applications [25]. We use BV programs with a secret string containing all 1s.
Pauli Evolution programs like UCCSD [26] are widely used to simulate the property of physics systems. To evaluate the averaged performance, we consider 10 random Pauli strings in the evolution where each Pauli string interaction involve at most $j$ qubits ($j$ is randomly chosen between 2 and $n$), for a $n$-qubit Pauli evolution circuit.
Hardware Modeling: When evaluating the design options of the quantum instruction set, we follow the established experimentally realistic hardware model for existing superconducting processors [3, 4]. We summarize some key points as follows:
  1. 1.
    For device Hamiltonian consisting of XX interactions between qubits [3], the latency $\tau$ of $D_{a},D_{b}$ and $D_{c}$ gates is proportional to the $L1$ norm of their Cartan coordinates (i.e., $\theta_{x}+\theta_{y}+|\theta_{z}|$);
  2. 2.
    For device Hamiltonian consisting of both XX and YY interactions [4] between qubits, the latency $\tau$ of $D_{a}$ and $D_{b}$ gates is proportional to the $L\infty$ norm of their Cartan coordinates (i.e., the value of $\theta_{x}$);
  3. 3.
    The error rate of $D_{a},D_{b}$ and $D_{c}$ gates is proportional to their latency $\tau$ [3, 4], and there is an experimentally-fitted formula: $p_{e}=m\tau+b$, $\frac{\pi}{4}\cdot m\approx 5.76\times 10^{-3}$, $b\approx 1.909\times 10^{-3}$.
(a) For QCCA with $D_{a}$
(b) For QCCA with $D_{b}$
(c) For QCCA with $D_{c}$
Figure 2: Comparing the decomposition cost by QCCA to that by NuOp. Both NuOp and QCCA adopt the same 2q basis gate in compilation. As for the X axis, for (a), x axis value $k$ means to use $D_{a}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k}$; for (b), x axis value $k$ means to use $D_{b}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k},\theta_{y}=\frac{\pi}{8k}$; for (c), x axis value $k$ means to use $D_{c}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k},\theta_{y}=\frac{\pi}{8k},\theta_{z}=\frac{\pi}{16k}$.
V-B Comparing to the baseline
Figure 3: Comparing QCCA to XXDecomposer (can only works with $D_{a}$ type gates). The y axis is $\log_{2}t_{XXDecomposer}/t_{QCCA}$, where $t_{XXDecomposer}$ and $t_{QCCA}$ are the compilation time by XXDecomposer and QCCA, respectively. As for the X axis in (a)(b)(c)(d), x axis value $k$ means to use $D_{a}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k}$.
The comparison of the proposed QCCA framework and the baseline NuOp in terms of compilation time overhead and decomposition cost are shown in Figure 1 and 2 respectively. Note that for the compilation by QCCA with $D_{c}$, we adopt the first method in Section IV-C which is purely analytical and does not need to solve Equation IV.9.
Compilation time comparison between QCCA and NuOp: As shown in Figure 1, the compilation time by the NuOp is thousands of times of that by QCCA, and the ratio grows bigger as the rotation angle becomes smaller. For example, considering the compilation with $D_{a}=e^{i\frac{\pi}{32}X_{0}X_{1}}$, the compilation time of the SWAP gate by NuOp is about $2.5*10^{5}$ times of the time by QCCA. The time reduction by QCCA is due to its analytical nature. For the NuOp, as the rotation angle in $D_{a},D_{b}$ or $D_{c}$ becomes smaller, a circuit with more layers of 2q gates is needed in NuOp to approximate the target unitary, according to Theorem III.3. Further, NuOp will take $O(36n^{2})$ ($n$ is the circuit layer count) time/resource to compute and store the gradient for the numerical search engine, so that to finally reach the compilation result. Instead, the compilation time/resource by QCCA is constant across unitary operators to be decomposed.
It is important to have a small compilation time for 2q quantum operators to make the compiler scalable for practical quantum computing. Many real-world quantum applications, such as Grover’s algorithm [27], often consists of millions of two-qubit quantum operators. With QCCA, $10^{6}$ two-qubit quantum operators can be processed within 20 minutes, no matter how small the rotation angle in the basis 2q gate is. Instead, for NuOp, even if the basis 2q gate is just the CX gate, compiling $10^{6}$ two-qubit quantum operators will take about 5 days. This huge amount of compilation time will hurt the advantage of quantum processing.
Decomposition cost comparison between QCCA and NuOp: As shown in Figure 2, the decomposition cost by QCCA is comparable to that by the NuOp, when using $D_{a}$ or $D_{b}$. The averaged decomposition cost difference between NuOp and QCCA is about 1 with $D_{a}$ and is about 0 with $D_{b}$, for the tested benchmark programs. This demonstrates that for $D_{a}$ and $D_{b}$, it is possible to approximate the optimal synthesis (i.e., NuOp) with the analytical decomposition technique. The result also demonstrates that $D_{b}$ may be a more efficient basis gate than $D_{a}$ (when $\theta_{x}$ is the same), as $D_{b}$ leads to lower decomposition cost for test programs. The more optimization opportunities by $D_{b}$ discussed in Section IV-B also makes it easier to generate decomposition as efficient as the optimal synthesis.
However, for the case with $D_{c}$, QCCA is not as efficient as NuOp. This is due to the difficulty in solving Equation IV.9 analytically. The indirect synthesis with $D_{c}$ in Section IV-C uses two $D_{c}$ to produce the gate of the $D_{a}$ type, making the decomposition cost with $D_{c}$ at least the multiple of 4, missing the optimization opportunity for the quantum operator whose compiled circuit indeed only requires $24k+$ 2 ($k\in\mathbb{Z}$) invocations of $D_{c}$. This demonstrates that it is necessary to take thousands of times more compilation time if we want to approximate the optimal synthesis with $D_{c}$. On the other hand, $D_{c}$ is naturally less efficient than $D_{b}$. No matter for NuOp or QCCA, the decomposition cost with $D_{c}$ is on average slightly larger than that by $D_{b}$ when the x axis value is the same, as shown in Figure 2.
Overall, with $D_{a}$ or $D_{b}$, the analytical decomposition by QCCA can shorten the compilation time by thousands of times when compared to the numerical decomposition by NuOp, at the cost of slightly increasing the decomposition cost.
Comparing QCCA to XXDecomposer: The comparison of the proposed QCCA framework and the baseline XXDecomposer in terms of compilation time overhead and decomposition cost are shown in Figure 3(a)(b)(c)(d), respectively. We can see that QCCA still achieves significant speedup (up to 206.4x for the tested configurations) comparing to XXDecomposer and this speedup grows as the rotation angle in $D_{a}$ become smaller. QCCA compiles all 2q operators globally (i.e., with a fixed compilation time), while XXDecomposer adopts a path-like searching for candidate circuits. Though XXDecomposer finds the candidate circuit analytically in each step of searching, the increased search step induced by a smaller rotation angle in $D_{a}$ still hurts the scalability of XXDecomposer.
As for the decomposition cost, QCCA achieves the same decomposition cost as XXDecomposer for the CX, SWAP and CRZ (controlled z-rotation) gates. For the random unitary operator, the local optimality of XXDecomposer leads to overall slightly lower compilation cost than QCCA, with a cost difference about 1 averaged over all tested configurations. However, comparing to other 2q quantum operators, the CX, SWAP and CRZ gates are most commonly-used among popular quantum applications, e.g., QAOA, QFT, BV and Pauli evolution circuits. Moreover, QCCA can work with nonstandard 2q basis gates of $D_{b}$ and $D_{c}$ types. As a comparison, it still remains unclear how to generalize XXDecomposer to work with $D_{b}$ and $D_{c}$.
V-C Empirical Design of the Quantum Instruction Set
In this section, we evaluate various options of the quantum instruction set design with the proposed analytical framework. An efficient quantum instruction set can improve the fidelity of decomposed circuits and reduce the program latency, enhancing the quantum advantage.
Figure 4: Comparing $D_{a}$ to CX on compiling various quantum applications. For the X axis in (a)(b)(c)(d), x axis value $k$ means to use $D_{a}$ with the rotation angle $\theta_{x}=\frac{\pi}{4k}$. QFT-5 means a 5-qubit QFT circuit. Other legend labels are named in the same manner.
(Q1) Why using basis 2q gate with a small rotation angle?
As shown in Figure 4(a), using $D_{a}$ with some small rotation angles (e.g., $\theta_{x}=\frac{\pi}{16}$) can improve the fidelity of QFT and QAOA circuits. This is because there are many small rotation angles in the CRZ gates of the QFT circuit and in the ZZ interaction gate of QAOA circuits. We may only use two invocations of $D_{a}$ to implement those gates. Even with the CX gate, we still need two invocations of CX gates to implement those gates. According to the experimentally-realistic error model [3, 4], the error rate of the CX gate is far larger than that of the $D_{a}$ gate. Thus, for circuits consisting of many 2q operators that has a small Cartan coordinate, using $D_{a}$ gate with a small rotation angle will provide considerable improvement for the fidelity of the decomposed circuit. We also points out that, the fidelity improvement is bounded by the amount rather than the percentage of 2q operators with a small Cartan coordinate, due to the fact that the circuit fidelity is computed as the product of individual gate fidelity. Thus, the fidelity improvement becomes larger when the QFT and QAOA program involve more qubits, as shown in Figure 4(a).
Likewise, the latency reduction induced by using $D_{a}$ with a small $\theta_{x}$ (see Figure 4(b)) is due to the same reason. However, this latency reduction is bounded by the percentage of 2q operators with a small Cartan coordinate, existing in QFT and QAOA circuits. This means the latency reduction induced by using a even smaller rotation angle in $D_{a}$ will gradually converge to some fixed value, as supported by the tendency shown in Figure 4(b). On the other hand, using $D_{a}$ with a too small rotation angle will hurt the circuit fidelity, as shown in Figure 4(a). Thus, to find the $D_{a}$ which maximally boosts the circuit performance, we need to make a good balance between the fidelity improvement and the latency reduction.
Compared to other state-of-the-art tools, QCCA can help finding the optimal rotation angle for $D_{a}$, $D_{b}$ or $D_{c}$ much more efficiently due to its extremely low time and resource overhead. QCCA can decompose the quantum program with $10^{3}$ different 2q operators and thus obtaining its fidelity and latency within one second, making QCCA useful not only for the quantum instruction set design in the near term, but also for future large-scale quantum applications.
(Q2) Will having mixed 2q basis gates in the instruction set be more efficient?
By comparing Figure 3(c) to (a), we can see that, for different benchmark programs, the rotation angle of $D_{a}$ that leads to the best circuit fidelity is also different. For the BV and Pauli evolution benchmark, most 2q operators are just the CX gate, and the fidelity of implemented CX with $D_{a}$ ($\theta_{x}<\frac{\pi}{4}$) is lower than the directly implemented CX on hardware. In this case, the fidelity improvement will decrease as the qubit count of the circuit increases or as $\theta_{x}$ decreases.
Therefore, having $D_{a}$ with different rotation angles in the instruction set is important for always obtaining circuit fidelity improvement over various quantum applications. For the tested eight benchmark programs, we prefer the mixed instruction set $\{CX,D_{a}(\frac{\pi}{16})\}$ if putting more emphasis on the fidelity improvement, or $\{CX,D_{a}(\frac{\pi}{32})\}$ if putting more emphasis on the latency reduction. This design also works for the $D_{b}$ gate since for those programs, the required number of invocations is the same for $D_{a}$ and $D_{b}$ if their $\theta_{x}$ is the same.
In the general case, hardware engineers can use QCCA to test more benchmark programs, and select appropriate rotation angles for 2q basis gates in an application-aware way demonstrated above. Researchers [2, 4] have proposed various calibration methods for nonstandard 2q basis gates. For a mixed instruction set $G$ with $m$ different basis gates, its calibration overhead is proportional to $m$. With QCCA, it is easy to take the calibration cost into the mixed instruction set design, only needing to add a term related to the calibration cost in the objective function that measures the efficiency of the instruction set $G$.
(Q3) Which type of gates is most efficient? $D_{a}$, $D_{b}$, or $D_{c}$?
We rule out $D_{c}$ for the following reasons. Firstly, the decomposition cost by $D_{b}$ is comparable to that by $D_{c}$ if their $\theta_{x}$ and $\theta_{y}$ are the same, as shown in Figure 2(b)(c). Secondly, $D_{c}$ gates are hard to prepare in existing quantum platforms [20, 5, 4], whose device Hamiltonian either only consists of XX interactions or XX+YY interactions. We must use $D_{a}$ or $D_{b}$ to implement $D_{c}$ in an indirect manner. Last but not least, as shown in Figure 2(c), when using $D_{c}$, it is hard to approach the optimal compilation with analytical schemes. This fact will cause great difficulty in designing a near-optimal and scalable compiler based on $D_{c}$.
Whether $D_{a}$ or $D_{b}$ is more efficient depends on the hardware platform. For the quantum processor based on the XX coupling between qubits, the instruction combination $G_{1}=\{D_{a}(\theta_{x}),D_{a}(\theta_{y})\}$ is more efficient than $G_{2}=\{D_{b}(\theta_{x},\theta_{y})\}$ since the gate $D_{b}(\theta_{x},\theta_{y})$ is implemented by $D_{a}(\theta_{x})S_{0,1}D_{a}(\theta_{y})S^{\dagger}_{0,1}$ in such a platform. Any fidelity improvement and latency reduction achieved by $G_{2}$ can thus also be achieved by $G_{1}$. On the other hand, $G_{1}$ can achieve efficient decomposition that cannot be achieved by $G_{2}$. For example, $e^{i(\theta_{x}+\theta_{y})X_{0}X_{1}}$ can be achieved by $D_{a}(\theta_{x})D_{a}(\theta_{y})$ and $D_{b}Y_{1}(\beta)D_{b}^{\dagger}$. It is obvious that the former circuit has a smaller latency and a higher fidelity.
For the hardware platform based on the XX+YY coupling between qubits, we argue $D_{b}$ is more efficient. In such a platform, for a fixed $\theta_{x}$, $D_{b}(\theta_{x},\theta_{y})$ has a similar latency and fidelity as $D_{a}(\theta_{x})=D_{b}(\theta_{x},0)$, according to the established hardware model [4]. While the performance of such $D_{b}$ and $D_{a}$ are almost identical for benchmark programs whose 2q operators are locally equivalent to XX interactions, the extra $\theta_{y}$ in $D_{b}$ can largely reduce the decomposition cost of the SWAP gate (see Figure 2(a)(b)), which appears frequently (up to 4x the count of other 2q operators [28]) in qubit-connectivity-limited quantum processors. We can further find the optimal $\theta_{y}$ for the fixed $\theta_{x}$ by examining the decomposition cost with QCCA. For example, for $\theta_{x}=\frac{\pi}{4}$, $\theta_{y}=\frac{\pi}{8}$ can achieve the best decomposition cost for both random unitary operators and the SWAP gate; for $\theta_{x}=\frac{\pi}{8}$, $\theta_{y}=\frac{\pi}{8}$ is the best choice.
The discussion above also invalidates the option of having both $D_{a}$ and $D_{b}$ in the quantum instruction set. This is because the most efficient type of gates are determined by the type of the qubit coupling. We may have both $D_{a}$ and $D_{b}$ in the instruction set only when the device Hamiltonian has mixed types of coupling between different qubit pairs.
VI Related Work
Works that directly relate to our paper target quantum compilation and quantum instruction design.
For quantum compilation works, most existing compilers [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 29, 20, 30, 31, 32, 33, 34, 5] assume using a few standard two-qubit gates like the CX/CZ and iSWAP gate in the compilation. Those compilers cannot work with an arbitrarily entangling two-qubit gate, let alone evaluate the pros and cons of different two-qubit basis gates.
On the other hand, researchers have tried to extend the border of the quantum instruction set architecture and design new quantum instructions, e.g., the continuous fsim gate family [35, 36], XY gate family [37], and CRISC gate family [4]. However, those works only provide analytical decomposition schemes for a few fixed two-qubit gates, like the $\sqrt{iSWAP}$ gate[38]. They do not support analytical decomposition with arbitrary 2q entangling gates. Lao et al.[1] and Lin et al.[2] propose to use the numerical compilation method so as to support all kinds of two-qubit basis gates. However, numerical methods often suffer from heavy time complexity, hindering wider exploration of the quantum instruction set design space. Peterson et al.[3] proposes a locally analytical compiler for all gates in the XX gate family. However, their tool cannot work with 2q basis gates of the $D_{b}$ and $D_{c}$ type.
VII Discussion and Conclusion
This paper proposes an analytical framework QCCA to analyze the Cartan coordinate of two-qubit operators and proposes a series of constructions to convert between 2q operators based on the feature of their Cartan coordinates. This work further proposes a compiler that significantly reduces the conversion cost between two-qubit operators and demonstrates comparable overall decomposition results with the numerically optimal compiler. Compared to the state-of-the-art, QCCA reduces the time overhead of quantum instruction set design by several orders of magnitude. With QCCA, this work evaluates and proposes feasible designs for the quantum instruction set, further enhancing quantum advantage. While this work establishes comprehensive analytical compilation with nonstandard 2q basis gates, there is still much space left for improvements.
(1) Better analytical compilation with $D_{c}$: This paper establishes the analytical compilation with $D_{c}$ by converting gates of the $D_{c}$ gate type into those off the $D_{a}$ type. Such conversion itself comes with a cost. Moreover, compared to $D_{a}$, one invocation of $D_{c}$ can simultaneously affect all Cartan coordinate directions of the resulting unitary. How to efficiently utilize this feature of $D_{c}$ still remains an open question. While it is hard to derive a global analytical solution to the Equation IV.9, it might be possible to design a local analytical scheme as the one presented in [3].
(2) Better triangle inequality for two-qubit operators: For a more general triangle inequality than the one presented in Theorem III.3, we may want to know if the following conjecture is true or not:
Conjecture.
For any $U\in SU(4)$, if $U$ can be implemented with n $D_{a}$ (or $D_{b}$ or $D_{c}$), then $k_{t}(U)\leq n\theta_{x}$ (or $\leq n(\theta_{x}+\theta_{y})$ or $\leq n(\theta_{x}+\theta_{y}+|\theta_{z}|)$).
A more accurate estimate of the decomposition cost can guide us to find a more optimal decomposition scheme and a faster evaluation of the effectiveness of a nonstandard 2q gate.
(3) More operation construction with nonstandard gates: This work mainly discusses how to use nonstandard 2q gates to construct other 2q operators. Another interesting direction is to explore how to use $D_{a}$ (or $D_{b}$ or $D_{c}$) to directly synthesize n-qubit quantum operations, such as the Pauli string evolution, without using the CX-based indirect synthesis. This may help us gain benefits with the nonstandard 2q gate that has a small Cartan coordinate, even if the benchmark program is full of CX-like gates.
Appendix
We illustrate the proof of Theorem III.3 as follows. The proof is based on the KAK decomposition, of which the details can be founded in [39]. For a given $U\in SU(4)$, consider
$M^{\dagger}UM=U_{r}+iU_{i}=Q_{L}(D_{r}+iD_{i})Q_{R}^{T},$
where $Q_{L},Q_{R}$ are orthogonal matrices, $D_{r},D_{i}$ are diagonal matrices, $M=\frac{1}{\sqrt{2}}\begin{pmatrix}1&0&0&i\\ 0&i&1&0\\ 0&i&-1&0\\ 1&0&0&-i\end{pmatrix},(D_{r}+iD_{i})=exp(i\,\text{diag}(\eta_{0},\eta_{1},\eta_{2},\eta_{3})),$ with $\eta_{i}$ determined by the global phase $g\in\{0,\frac{\pi}{2}\}$ and the Cartan coordinate ($k_{x},k_{y},k_{z}$), $\frac{\pi}{4}\geq k_{x}\geq k_{y}\geq|k_{z}|\geq 0$:
$\begin{pmatrix}\eta_{0}\\ \eta_{1}\\ \eta_{2}\\ \eta_{3}\\ \end{pmatrix}=\begin{pmatrix}+1&+1&-1&+1\\ +1&+1&+1&-1\\ +1&-1&-1&-1\\ +1&-1&+1&+1\end{pmatrix}\begin{pmatrix}g\\ k_{x}\\ k_{y}\\ k_{z}\\ \end{pmatrix}$ (VII.1)
Without ambiguity, here $\sigma_{i}(U)$ refers to the $i$-th maximal singular value of $U$.
Proposition VII.1.
If $g=0$, then the singular value of $U_{i}$ is:
$\begin{pmatrix}\sigma_{1}(U_{i})\\ \sigma_{2}(U_{i})\\ \sigma_{3}(U_{i})\\ \sigma_{4}(U_{i})\end{pmatrix}=\begin{pmatrix}\sin(k_{x}+k_{y}+k^{+}_{z})\\ \sin(k_{x}+k_{y}-k^{+}_{z})\\ \sin(k_{x}-k_{y}+k^{+}_{z})\\ \sin|-k_{x}+k_{y}+k^{+}_{z}|\end{pmatrix},$ (VII.2)
where $k_{z}^{+}=|k_{z}|$. If $g=\frac{\pi}{2}$, we only need to replace $U_{i}$ in Equation VII.2 with $U_{r}$.
Proof.
Firstly, $|D_{i}|$ contains all singular values of $U_{i}$ regardless the ordering. For example, $|\sin\eta_{2}|$ is just $\sin(k_{x}+k_{y}+k_{z})$ if $k_{z}\geq 0$. As for the ordering of those singular values:
(1) $\sin(k_{x}+k_{y}+k^{+}_{z})\geq\sin(k_{x}+k_{y}-k^{+}_{z})\Longleftrightarrow k_{x}+k_{y}+k^{+}_{z}\leq\pi-(k_{x}+k_{y}-k^{+}_{z})\Longleftrightarrow k_{x}+k_{y}\leq\frac{\pi}{2}$;
(2) $\sin(k_{x}+k_{y}-k^{+}_{z})\geq\sin(k_{x}-k_{y}+k^{+}_{z})\Longleftrightarrow k_{y}\geq k_{z}^{+}$;
(3) $\sin(k_{x}-k_{y}+k^{+}_{z})\geq\sin|-k_{x}+k_{y}+k^{+}_{z}|\Longleftrightarrow k_{x}-k_{y}+k^{+}_{z}\geq|-k_{x}+k_{y}+k^{+}_{z}|$
If $k_{y}+k^{+}_{z}\geq k_{x}$, the inequality is equivalent to $k_{x}\geq k_{y}$;
If $k_{y}+k^{+}_{z}\leq k_{x}$, the inequality is equivalent to $k_{z}^{+}\geq 0$;
Therefore, the singular values of $U_{i}$ is just as shown in the proposition. The case for $U_{r}$ can be computed similarly. ∎
We define $k_{t}(U)=(k_{x}+k_{y}+k_{z}^{+})(U)$, $R(\beta)=\min(\beta,\pi-\beta)$, $\theta=\theta_{x}$ (of $D_{a}$). We assume $0<\theta<\frac{\pi}{4}$.
Proposition VII.2 (Triangle inequality).
For $U\in SU(4)$, if $R(k_{t}(U))\leq\frac{\pi}{4}$, $|\cos k_{t}(UD_{a})|\leq\cos(R(k_{t}(U))-\theta)$.
Proof. Let $U_{1}=UD_{a}$, then
$U_{1r}+iU_{1i}$ $=M^{\dagger}U_{1}M=M^{\dagger}UMe^{i\theta Z_{0}}$
$=(\cos\theta U_{r}-\sin\theta U_{i}Z_{0})+i(\cos\theta U_{i}+\sin\theta U_{r}Z_{0})$
Since $Z_{0}\in SO(4)$, the singular values of $U_{i}Z_{0}$ should be as the same as $U_{i}$. Then according to Weyl’s Inequality [40],
$\sigma_{4}(U_{1r})\leq\cos\theta\sigma_{4}(U_{r})+\sin\theta\sigma_{1}(U_{i}),$ (VII.3)
$\sigma_{4}(U_{1i})\leq\sin\theta\sigma_{4}(U_{i})+\cos\theta\sigma_{1}(U_{r}).$ (VII.4)
$\sigma_{4}(U_{1r})\leq\cos\theta\sigma_{1}(U_{r})+\sin\theta\sigma_{4}(U_{i}),$ (VII.5)
$\sigma_{4}(U_{1i})\leq\sin\theta\sigma_{1}(U_{r})+\cos\theta\sigma_{4}(U_{i}),$ (VII.6)
Now, we do case by case study regarding the global phase $g_{U_{1}}$ and $g_{U}$ of $U_{1}$ and $U$, respectively.
(i) If $g_{U_{1}}=0$ and $g_{U}=0$, then by Proposition VII.1, Equation VII.3 is equivalent to
$|\cos(k_{t}(U_{1}))|\leq$ $\cos\theta|\cos k_{t}(U)|+\sin\theta\sin(k_{t}(U)),$
$=\cos(R(k_{t}(U))-\theta)$ (VII.7)
This completes the proof.
(ii) If $g_{U_{1}}=\frac{\pi}{2}$ and $g_{U}=0$, then by Proposition VII.1, Equation VII.4 becomes
$|\cos(k_{t}(U_{1}))|$ $\leq\cos\theta sink_{t}(U)+\sin\theta|\cos(k_{t}(U))|$
$=\sin(R(k_{t}(U))+\theta)\leq\cos(R(k_{t}(U))-\theta).$
This completes the proof. The second inequality is due to $R(k_{t}(U))\leq\frac{\pi}{4}$.
Other cases, such as $(g_{U_{1}},g_{U})=(0,\frac{\pi}{2})$ and $(g_{U_{1}},g_{U})=(\frac{\pi}{2},\frac{\pi}{2})$ can be discussed similarly.
Now we start to prove Theorem III.3.
Proof.
The theorem can be proved by contradiction:
Assume $sin\,k_{t}(U)>\sin N\theta$. This means $R(k_{t}(U))>N\theta$. Now consider $U_{1}=U(a_{1}\otimes b_{1})D_{a}^{\dagger}(c_{1}\otimes d_{1})$, where $a_{1},b_{1},c_{1},d_{1}\in SU(2)$. Note that single-qubit gates do not affect the Cartan coordinate. Then we discuss $R(k_{t}(U_{1}))$ case by case:
(i) if $R(k_{t}(U))\leq\frac{\pi}{4}$, $|\cos k_{t}(U_{1})|\leq\cos(R(k_{t}(U))-\theta)$, according to Proposition VII.2. This indicates $R(k_{t}(U_{1}))\geq R(k_{t}(U))-\theta$.
(ii) If $R(k_{t}(U))>\frac{\pi}{4}$ and $g_{U_{1}}=g_{U}$, then we still have $R(k_{t}(U_{1}))\geq R(k_{t}(U))-\theta$, following the proof of Proposition VII.2;
(iii) If $R(k_{t}(U))>\frac{\pi}{4}$ and $g_{U_{1}}\neq g_{U}$, following the proof of Proposition VII.2 (now using Equation VII.6 for the case (ii) in the proof of Proposition VII.2), we have $|\cos(k_{t}(U_{1}))|\leq\cos\theta\sin|-k_{x}+k_{y}+k_{z}^{+}|+\sin\theta\cos|-k_{x}+k_{y}+k_{z}^{+}|=\sin|-k_{x}+k_{y}+k_{z}^{+}|+\theta\leq\sin(\frac{\pi}{4}+\theta)=\cos(\frac{\pi}{4}-\theta)$. This leads to $R(k_{t}(U_{1}))\geq\frac{\pi}{4}-\theta$.
Overall, we have $R(k_{t}(U_{1}))\geq\min(\frac{\pi}{4},R(k_{t}(U)))-\theta>(N-1)\theta.$ Then for $U_{N}=U_{N-1}(a_{N-1}\otimes b_{N-1})D_{a}^{\dagger}(c_{N-1}\otimes d_{N-1})$, with recursion, $R(k_{t}(U_{N}))>(N-N)\theta=0$.
However, by adjusting single-qubit gates $\{a_{i},b_{i},c_{i},d_{i}\}$, the generated set of $U_{N}$ should contain the identity gate. In this case, $00=R(k_{t}(I))=R(k_{t}(U_{N}))>$ 0, causing the contradiction. Thus, $\sin k_{t}(U)\leq\sin N\theta$. ∎